OpenStudy (elvisiscool69):

what is f(-3) if f(x)=|2x+1|+5? 1 point a.0 b.10 c.12 d.-10

OpenStudy (elvisiscool69):

@jigglypuff314

OpenStudy (ayeedomo):

Plug in the (-3) in place of all x values

OpenStudy (ayeedomo):

f(-3)=|2(-3)+1|+5

OpenStudy (elvisiscool69):

is it 10?

OpenStudy (elvisiscool69):

@ayeedomo

OpenStudy (elvisiscool69):

or -10

OpenStudy (ayeedomo):

hang on for a moment I'll check :)

OpenStudy (elvisiscool69):

i could be wrong tho

OpenStudy (ayeedomo):

f(-3)=|2(-3)+1|+5 f(-3) -5 = -6 + 1 f(-3) -5 = -5 f(-3) = 0

OpenStudy (ayeedomo):

There's a small chance I could be wrong, because I'm not sure whether the absolute value bars can be dropped after subtracting the 5.

OpenStudy (elvisiscool69):

i could be wrong too! lol omg so what do we do???!!

OpenStudy (anonymous):

10 is right 2(-3) is equal to -6 2(-3)+1 is equal to -5 |2(-3)+1| is equal to 5 so 5+5 is 10

OpenStudy (ayeedomo):

I'd go with 10 then, pay no attention to my incorrect calculations XD Let me know if it turns out to be the correct answer

OpenStudy (elvisiscool69):

lol yeah man i'll let you kknow for sure

OpenStudy (elvisiscool69):

it is 10!

OpenStudy (ayeedomo):

XD